# CODI: Compressing Chain-of-Thought into Continuous Space via Self-Distillation

### v1

@@ Abstract
Chain-of-Thought (CoT) enhances Large Language Models (LLMs) by enabling step-by-step reasoning in natural language. However, the language space may be suboptimal for reasoning. While implicit CoT methods attempt to enable reasoning without explicit CoT tokens, they have consistently lagged behind explicit CoT method in task performance. We propose CODI (Continuous Chain-of-Thought via Self-Distillation), a novel framework that distills CoT into a continuous space, where a shared model acts as both teacher and student, jointly learning explicit and implicit CoT while aligning their hidden activation on the token generating the final answer. CODI is the first implicit CoT method to match explicit CoT’s performance on GSM8k while achieving 3.1× compression, surpassing the previous state-of-the-art by 28.2% in accuracy. Furthermore, CODI demonstrates scalability, robustness, and generalizability to more complex CoT datasets. Additionally, CODI retains interpretability by decoding its continuous thoughts, making its reasoning process transparent. Our findings establish implicit CoT as not only a more efficient but a powerful alternative to explicit CoT.
@@
Zhenyi Shen1, Hanqi Yan1, Linhai Zhang1, Zhanghao Hu1, Yali Du1,2, Yulan He1,2 1King’s College London           2The Alan Turing Institute {zhenyi.shen, yanqi.yan, linhai.zhang, zhanghao.hu}@kcl.ac.uk {yali.du, yulan.he}@kcl.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable reasoning capabilities OpenAI (2024); Anthropic (2024); Google (2024), with Chain-of-Thought (CoT) Wei et al. (2022) emerging as a key technique for enabling step-by-step reasoning via natural language rationales.
However, neuroscientific studies Amalric and Dehaene (2016, 2019) show that human mathematical reasoning does not primarily involve language processing areas in brains, suggesting that natural language may not be the most effective medium for reasoning. Furthermore, Lin et al. (2025) highlights that LLMs often depend heavily on specific tokens during reasoning, which can lead to errors despite the tokens being commonsensically valid. Replacing these tokens with alternatives has been shown to improve performance, underscoring LLMs’ sensitivity to linguistic features in natural language rationales. These insights motivate a shift from natural language CoT representations (Explicit CoT and Discrete Tokens) to dense, continuous representations (Implicit CoT and Continuous Thoughts) that may better align with the compact and abstract nature of reasoning.
Figure 1: Illustration of the training process of different reasoning approaches. CoT-SFT is standard CoT finetuning. Coconut learns implicit CoT by curriculum learning. CODI learns implicit CoT by self-distillation.
Implicit CoT requires two major design considerations: the forward function and the training objective. Various forward functions have been explored, including removing all reasoning tokens Deng et al. (2023, 2024), adding fixed learning tokens Pfau et al. (2024); Goyal et al. (2024), performing autoregression by connecting the last hidden activation to the next input embedding Hao et al. (2024). The autoregression approach appears to perform the best, likely because it increases the effective computational depth. However, the key design challenge for implicit CoT lies in the training objective. While explicit CoT learns reasoning through language modeling over annotated CoT tokens, implicit CoT cannot rely on this standard language modeling approach, as it must avoid generating explicit CoT tokens by definition.
To address this challenge, Coconut Hao et al. (2024), the state-of-the-art method, adopts a curriculum learning strategy initially introduced by Deng et al. (2024) as illustrated in Figure 1. It gradually replaces the initial CoT tokens with continuous thoughts while maintaining the language modeling objective on the remaining CoT tokens. In this way, the language modeling loss encourages the continuous thoughts to behave like the removed CoT tokens, and at the final stage of learning, all CoT tokens are replaced with continuous thoughts, achieving full implicit CoT. However, while Coconut outperforms the No-CoT baseline (which entirely omits CoTs), it still lags behind CoT-SFT by 20% on GSM8k, exposing a key limitation in the implicit CoT paradigm. We believe these limitations stem from the curriculum learning strategy itself—the multi-stage process delays the acquisition of a complete discrete CoT. If the model fails to generate the correct continuous thought at any stage due to forgetting or incomplete learning, errors propagate through subsequent stages, limiting overall performance.
We propose CODI (Continuous Chain-of-Thought via Self-Distillation), a novel framework that distills explicit CoT into implicit CoT by aligning the hidden activation of the token responsible for generating the final answer as illustrated in Figure 1. CODI reframes implicit CoT learning as a self-distillation task Wang et al. (2023); Gou et al. (2021), where the same model serves as both teacher and student. Unlike conventional self-distillation having the two roles of equal capability, CODI enhances the teacher’s knowledge by providing them distinct input contexts: the teacher learns from the groundtruth CoT and final answer using a language modeling objective, while the student generates continuous thoughts before predicting the final answer–our target task. Distillation occurs at the token preceding the final answer, which Orgad et al. (2025) identify as encoding crucial reasoning information. Since we can formally show that CoT influences the hidden activation of this token only by a shift (Section 3.3), CODI enforces alignment between the teacher and student by minimizing their hidden activation differences using an L1 distance loss, effectively injecting explicit CoT supervision into implicit CoT generation. This single-step distillation in feature space mitigates the forgetting issues inherent in curriculum learning, enabling more effective implicit CoT training.
@@
  • We demonstrate the effectiveness of distilling knowledge from explicit CoT (teacher) to implicit CoT (student) by aligning the hidden activation of a single token, simplifying the distillation process and improving efficiency.
  • Extensive experiments show that CODI is robust, scalable, and generalizable to more complex CoT datasets. Additionally, CODI maintains interpretability, making its reasoning process transparent.
Figure 2: CODI enables the model to generate continuous CoTs by jointly training a student and teacher task within a shared LLM, distilling knowledge from the teacher to the student. The Student task (left) generates the answer by autoregressively decoding continuous thoughts, while the Teacher task (right) generates the answer using the groundtruth CoT via teacher forcing. Both tasks learn the generated texts via cross-entropy loss ($\mathcal{L}_{\text{student}}$ and $\mathcal{L}_{\text{teacher}}$), and share the same LLM. Knowledge distillation is achieved by applying $\mathcal{L}_{\text{KD}}$ (L1 loss) between student and teacher hidden activation across all layers ($\textbf{h}_{\text{student}}$ and $\textbf{h}_{\text{teacher}}$).
@@ 2 Related Work
Implicit CoT methods aim to enhance reasoning without verbalizing intermediate steps as in CoT, thereby accelerating inference speed. Theoretical work Strobl et al. (2024); Merrill and Sabharwal (2024) establishes that additional computational tokens enhance transformers’ reasoning capacity. Empirical studies Pfau et al. (2024); Goyal et al. (2024) validate these insights by training LLMs with extra dummy tokens before answering though in a limited scale and effect. Recent efforts Deng et al. (2023, 2024) distills CoT reasoning by fine-tuning. They improve over the No-CoT baseline, but fall behind CoT finetuning possibly due to discarding all intermediate tokens. Addressing this, Coconut Hao et al. (2024) reintroduces intermediate reasoning tokens via autoregressive hidden state propagation, combining curriculum learning from Deng et al. (2024). While this achieves some improvement over Deng et al. (2024), Coconut still lags behind explicit CoT, which we attribute to forgetting in curriculum learning. CODI replaces curriculum learning with a novel self-distillation framework, enabling a single-step learning process that avoids forgetting issues. Our work is also inspired by in-context compression Ge et al. (2024); Li et al. (2024b), though our work is compressing the generation instead of the existing contexts.
@@ 3 CODI: Continuous Chain-of-Thought via Self Distillation
Unlike traditional CoT reasoning, CODI bypasses autoregression in the vocabulary space, and directly connects the last hidden representation to the subsequent input. The key challenge in training such a model with continuous thoughts lies in designing an appropriate training objective. Conventional reasoning learning in explicit CoT fine-tuning relies on a language modeling objective over annotated CoT tokens, which inevitably leads to discrete CoT token generation—contradicting the definition of implicit CoT.
@@ 3.1 Overview
CODI addresses this by introducing a self-distillation framework (Figure 2) with two training tasks: a teacher task and a student task. The teacher task learns explicit CoT generation, while the student task learns implicit continuous CoT generation. Knowledge distillation is achieved by aligning the hidden activation of a key token from the teacher to the student via $\mathcal{L}_{KD}$. The overall training objective is a weighted sum of three losses, which will be detailed later:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters controlling the balance among the objectives..
A Python implementation of this framework is provided in Figure A3.
@@ 3.2 Student Task
The student task (Figure 2, left), the target task, generates continuous thoughts by autoregressively propagating the last hidden states and learns to generate the answer token using a cross-entropy loss:
@@
where $P$ is the probability distribution of the LLM, $y$ refers the answer label, $Q$ refers the question tokens, and $Z$ refers the continuous thoughts.
On its own, the model benefits only marginally from the additional computation Goyal et al. (2024) compared with the No-CoT scenario because there are no supervision for the continuous thoughts.
Additionally, CODI applies modifications exclusively to the student task. Two special tokens, bot and eot, mark the start and end of continuous reasoning, inspired by Hao et al. (2024). A two-layer MLP followed by layer normalization transforms the hidden representations of continuous thought tokens before feeding them into the next step.
3.3 Teacher Task
Unlike the student task, the teacher task (Figure 2, right) performs explicit CoT generation using a language modeling objective:
$\mathcal{L}_{\text{teacher}}=-\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}\mid y_{1:i-1},Q),$ (3)
where $y$ refers both the CoT and the answer labels, and $Q$ refers the question tokens.
The teacher task serves two key functions: (1) Reference Learning: By learning explicit CoTs, the teacher task equips the model with structured reasoning patterns, offering a foundational reference for the student task. (2) Latent Supervision: As the teacher has access to ground-truth CoT tokens, its hidden activation at the answer-generating token encapsulate essential reasoning information by attending to all preceding CoT tokens. In contrast, the student initially operates without such structured guidance. To address this disparity, CODI aligns the hidden activation of this token between the teacher and student across all layers using an L1 loss:
@@ 3.3 Teacher Task
where $M$ indicates the number of layers in the LLM, sg denotes stop gradient, and $\textbf{h}^{l}$ is the hidden activation of the LLM’s $l$-th layer.
The Distilled Token.
Rather than aligning with all tokens in the generated sentence, we select a distillation token for alignment. Inspired by the recent observations (Orgad et al., 2025) that the hidden activation of the token intermediately preceding the answer, i.e., the colon (“:”) in the answer prompt “The answer is:” (as shown in Figure 2), encodes far more information than output logits. We select this token’s hidden activation, h, for distillation. This selection can be further verified by the shift mechanism in in-context learning. Recent work Li et al. (2024a); Liu et al. (2023) demonstrates that in-context examples influence the final query token by shifting its hidden activation values. Extending this idea, we show that CoT tokens similarly induce a shift in hidden activation values of this target token, as formalized in Equation 5:
where q is the query of this target token, $\mathbf{h}^{l}_{\text{CoT}}$ is the hidden activation at layer $l$ with CoT (equivalent to $\mathbf{h}^{l}_{\text{teacher}}$), $\mathbf{h}^{l}_{\text{no-CoT}}$ is the corresponding activation without CoT, and the remaining term quantifies the shift introduced by the CoT rationale $R$.
This suggests that the target token’s hidden activation encode the influence of preceding reasoning steps, and $\mathbf{h}^{l}_{\text{student}}$ can learn this shift by minimizing a simple distance metric, such as L1 loss, with $\mathbf{h}^{l}_{\text{teacher}}$. A formal proof of this “CoT shift” phenomenon is provided in Appendix B.
@@ 3.4 Training and Inference
The continuous thoughts are generated dynamically during training, as they are not known beforehand. To achieve this, we decode them step by step, with a cache storing previous keys and values to maintain efficiency. When applying distance loss between two hidden activation, we observed a significant norm variations across layers Deng et al. (2023); Cheng and Durme (2024). To address this, we normalize each layer’s activation by dividing them by the standard deviation of the teacher’s hidden activation within the current batch.
For the distillation task, we employed the same model for the teacher task and the student task for two reasons: (1) Warm-up: When the teacher trains alongside the student, it creates a warm-up effect for $\mathcal{L}_{\text{KD}}$. Both components start from the same initialization point, diverge during training, and gradually converge as the student adapts. In contrast, a static pre-trained teacher initially presents an overly challenging objective, as its hidden states reflect fully developed reasoning patterns that the untrained student cannot immediately match. (2) Shared model representations: Using the same model mitigates alignment issues in hidden activation that arise when using separate models, enabling smoother and more effective information transfer between the teacher and student. The corresponding ablation studies, which validate these findings, are detailed in Table 3.
For training data, we exclude the final CoT step—the step responsible for generating the final answer—because including this step could allow the teacher’s hidden activation to take a shortcut. Specifically, the model might directly copy the result from the last CoT step to the token responsible for generating the exact answer token, bypassing the reasoning process. This behavior would undermine the quality of the target hidden activation, as they would no longer fully encode the reasoning patterns. The ablation results demonstrating the impact of this exclusion are presented in Table 3.
@@
Method In-Domain Out-of-Distribution
GSM8k GSM8k-NL SVAMP GSM-Hard MultiArith
GPT-2 Small
CoT-SFT 44.1% 34.8% 41.8% 9.8% 90.7%
No-CoT-SFT 19.1% 19.1% 16.4% 4.3% 41.1%
iCoT 30.1%* 3.2% 29.4% 5.7% 55.5%
Coconut 34.1%* -- -- -- --
CODI (Ours) 43.7% 35.3% 42.9% 9.9% 92.8%
LLaMA3.2-1b
CoT-SFT 61.6% 54.1% 66.7% 15.6% 99.3%
No-CoT-SFT 30.9% 30.9% 44.1% 7.1% 70.9%
iCoT 19.0% 15.2% 40.9% 4.4% 39.0%
CODI (Ours) 55.6% 49.7% 61.1% 12.8% 96.1%
Table 1: Results on four datasets: GSM8k, SVAMP, GSM-Hard, and MultiArith. GSM8k-NL indicates that the training data is GSM8k-Aug-NL, the verbose dataset, instead of GSM8k-Aug. Results marked with * are taken from the corresponding papers Deng et al. (2024); Hao et al. (2024). Coconut’s results are incomplete due to the unavailability of open-source code.

### v2

@@ Abstract
Chain-of-Thought (CoT) reasoning enhances Large Language Models (LLMs) by encouraging step-by-step reasoning in natural language. However, leveraging a latent continuous space for reasoning may offer benefits in terms of both efficiency and robustness. Prior implicit CoT methods attempt to bypass language completely by reasoning in continuous space but have consistently underperformed compared to the standard explicit CoT approach. We introduce CODI (Continuous Chain-of-Thought via Self-Distillation), a novel training framework that effectively compresses natural language CoT into continuous space. CODI jointly trains a teacher task (Explicit CoT) and a student task (Implicit CoT), distilling the reasoning ability from language into continuous space by aligning the hidden states of a designated token. Our experiments show that CODI is the first implicit CoT approach to match the performance of explicit CoT on GSM8k at the GPT-2 scale, achieving a 3.1x compression rate and outperforming the previous state-of-the-art by 28.2% in accuracy. CODI also demonstrates robustness, generalizable to complex datasets, and interpretability. These results validate that LLMs can reason effectively not only in natural language, but also in a latent continuous space.
@@
Zhenyi Shen1, Hanqi Yan1, Linhai Zhang1, Zhanghao Hu1, Yali Du1,2, Yulan He1,2 1King’s College London           2The Alan Turing Institute {zhenyi.shen, hanqi.yan, linhai.zhang, zhanghao.hu}@kcl.ac.uk {yali.du, yulan.he}@kcl.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable reasoning capabilities OpenAI (2024); Anthropic (2024); Google (2024), with Chain-of-Thought (CoT) Wei et al. (2022) emerging as a key technique for enabling step-by-step reasoning. The success of CoT can be explained as it allows human-like deliberate thinking when computing a sequence of reasoning tokens before deriving the final answer  (Kahneman, 2011).
However, conventional CoT-based methods only rely on natural language tokens as the medium for reasoning. While prior work on prompt learning Lester et al. (2021) has demonstrated that transforming discrete prompts into continuous representations can lead to efficient yet effective reasoning Li and Liang (2021). This motivates us to investigate if CoT reasoning can similarly benefit from continuous representations. Compared to natural language, reasoning in continuous space offers the following advantages. First, verbalizing the reasoning process can be inefficient, as many tokens are devoted to communication rather than computation Li et al. (2024b). Second, learning annotated CoTs token-by-token may cause models to overfit on superficial linguistic cues Lin et al. (2025). While continuous representations—without the need to mimic explicit targets—introduce a softer prior, which may lead to improved robustness.
Figure 1: Comparison of reasoning strategies. No-CoT-SFT: Train model on (Q,A) pairs via SFT. CoT-SFT: Train model on (Q, CoT, A) triples via SFT, i.e., with explicitly annotated CoT reasoning steps. Coconut: requires multi-stage training to progressively replace CoT tokens with continuous representations. CODI: achieves this in a single stage by compressing CoT tokens into continuous space via self-distillation.
An implicit CoT algorithm replaces natural language tokens with continuous representations for reasoning as shown in Figure 1 (left). To effectively learn these representations, Pfau et al. (2024); Goyal et al. (2024) pretrain the model with additional thinking tokens from scratch. More recently, the state-of-the-art method, Coconut Hao et al. (2024) adopts a curriculum learning strategy Deng et al. (2024) that gradually replaces the initial CoT tokens with continuous thoughts. This strategy encourages continuous thoughts to behave like the removed CoT tokens. Although Coconut has greatly improved upon earlier implicit CoT methods in terms of performance  (Goyal et al., 2024; Deng et al., 2024), it lags behind CoT-SFT by a large margin as shown in Figure 1 (right). We hypothesize that this performance gap is due to forgetting across stages in the curriculum learning process Vijjini et al. (2021). This prompts us to ask: Can implicit CoT methods achieve the reasoning capability comparable to CoT-SFT while maintaining their efficiency advantages?
To address this, we propose a novel training framework: CODI (Continuous Chain-of-Thought via Self Distillation). CODI enables implicit CoT learning in a single training step by leveraging self-distillation, thereby avoiding the forgetting issues inherent in curriculum learning. In doing so, it achieves performance comparable to CoT-SFT while being significantly more efficient. CODI enables implicit CoT reasoning through a joint learning setup involving a teacher task and a student task. The teacher learns from the annotated CoT tokens using a cross-entropy loss, while the student generates a small number of continuous thoughts before producing the final answer, representing implicit CoT reasoning. We do not constrain the student’s continuous thoughts to match any specific target. Instead, we transfer the teacher’s reasoning knowledge to the student through a form of representation alignment at the position of answer generation, where the essence of the reasoning process is captured Orgad et al. (2025). This allows the student to effectively mimic the teacher’s reasoning pattern in continuous space without rigid constraints. We refer to this mechanism as self-distillation Wang et al. (2023); Gou et al. (2021), emphasizing the model’s ability to distill one of its own behaviors into another.
@@
  • We demonstrate the effectiveness of distilling knowledge from explicit CoT to implicit CoT by aligning the hidden activations of a single token.
  • Extensive experiments show that CODI is robust, generalizable to complex CoT datasets, and offers a reasonable level of interpretability.
Figure 2: CODI enables the model to generate implicit continuous CoTs by jointly training a student task and a teacher task, and distills knowledge from the teacher to the student. The Student task (left) generates the answer by autoregressively decoding continuous thoughts starting from a learnable bot token, while the Teacher task (right) generates the answer using the groundtruth CoT via teacher forcing. Both tasks learn the generated texts via cross-entropy loss ($\mathcal{L}_{\text{student}}$ and $\mathcal{L}_{\text{teacher}}$), and share the same LLM. Knowledge distillation is achieved by applying $\mathcal{L}_{\text{KD}}$ (L1 loss) between student and teacher hidden activation across all layers ($\textbf{h}_{\text{student}}$ and $\textbf{h}_{\text{teacher}}$).
@@ 2 Related Work
Implicit CoT methods aim to enhance reasoning without verbalizing intermediate steps as in CoT, thereby accelerating inference speed. Theoretical work Strobl et al. (2024); Merrill and Sabharwal (2024) establishes that additional computational tokens enhance transformers’ reasoning capacity. Empirical studies Pfau et al. (2024); Goyal et al. (2024) validate these insights by training LLMs with extra dummy tokens before answering though in a limited scale and effect. Recent efforts Deng et al. (2023, 2024) distills CoT reasoning by fine-tuning. They improve over the No-CoT baseline, but fall behind CoT finetuning possibly due to discarding all intermediate tokens. Addressing this, Coconut Hao et al. (2024) reintroduces intermediate reasoning tokens via autoregressive hidden state propagation, combining curriculum learning from Deng et al. (2024). While this achieves some improvement over Deng et al. (2024), Coconut still lags behind explicit CoT, which we attribute to forgetting in curriculum learning. CODI replaces curriculum learning with a novel self-distillation framework, enabling a single-step learning process that avoids forgetting issues. Our work is also inspired by in-context compression Ge et al. (2024); Li et al. (2024c), though our work is compressing the generation instead of the existing contexts. Concurrent works Xu et al. (2025); Liu et al. (2024); Su et al. (2025) explore latent reasoning, but still rely on explicit CoT generation. Looped transformers Geiping et al. (2025a); Saunshi et al. (2025); Yu et al. (2025) also support latent reasoning, though they primarily vary in model depth without introducing. In contrast, CODI emphasizes increasing reasoning capability through additional tokens.
@@ 3 CODI: Continuous Chain-of-Thought via Self Distillation
Unlike traditional CoT reasoning, CODI bypasses autoregression in the vocabulary space, and directly connects the last hidden representation to the subsequent input. The key challenge in training such a model with continuous thoughts lies in designing an appropriate training objective. Conventional reasoning learning in explicit CoT fine-tuning relies on a cross-entropy loss over annotated CoT tokens, which inevitably leads to discrete CoT token generation—contradicting the definition of implicit CoT.
@@ 3.1 Overview
CODI addresses this challenge by introducing a self-distillation framework (Figure 2) with two training tasks: a teacher task and a student task. The teacher task learns explicit CoT reasoning, while the student task learns implicit CoT reasoning. Knowledge distillation is achieved by aligning the hidden activations of a key token from the teacher to the student via $\mathcal{L}_{KD}$. The overall training objective is a weighted sum of three losses:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters controlling the balance among the objectives. [^1] A Python implementation of this framework is provided in Figure A3.
3.2 Teacher Task
The teacher task (Figure 2, right) learns explicit CoT using a cross-entropy loss:
$\mathcal{L}_{\text{teacher}}=-\frac{1}{N}\sum_{i=1}^{N}\log P(r_{i}\mid r_{1:i-1},Q),$ (2)
where $P$ denotes the output probability distribution of the LLM, $Q$ represents the question tokens, and $r=\left[c,y\right]$ is the concatenated sequence of the CoT reasoning tokens $c$ and the final answer token $y$.
@@ 3.3 Student Task
The student task (Figure 2, left), which performs implicit CoT reasoning, generates continuous thoughts by autoregressively propagating the last hidden states. This process begins with a learnable <bot> (begin-of-thoughts) token and proceeds until a learnable <eot> (end-of-thoughts) token is reached. The model then learns the final answer from the <eot> token using a cross-entropy loss:
@@
where $y$ denotes the answer label, $Q$ the question tokens, and $Z$ the continuous thoughts.
Additionally, a two-layer MLP followed by layer normalization transforms the hidden representations of continuous thought tokens before feeding them into the next step for the purpose of better discriminating the latent space and the token space.
3.4 Self-Distillation
If the model learns only with the student task, it benefits only marginally from the additional computation Goyal et al. (2024) due to the absence of supervision for continuous thoughts.
Distillation in Feature Space.
To provide explicit supervision to guide continuous thoughts, we adopt a feature-level distillation strategy. Recent work Li et al. (2024a); Liu et al. (2023) demonstrates that in-context examples influence the final query token by shifting its hidden activation values. Extending this idea, we show that CoT tokens similarly induce a shift in hidden activation values of a query token (can be a probing token like "Answer") compared to a sequence without CoT, as formalized in Equation 4:
where q is the query token, $\mathbf{h}^{l}_{\text{CoT}}$ is the hidden activations at layer $l$ with CoT, $\mathbf{h}^{l}_{\text{no-CoT}}$ is the corresponding activation without CoT, and the remaining term quantifies the shift introduced by the CoT rationale $R$. A formal proof of this “CoT shift” phenomenon is provided in Appendix B.
This decomposition suggests that the key information from CoT reasoning accessible to the query token is embedded in the shift term $f(\cdot)$. Therefore, by encouraging the student’s hidden activations $\mathbf{h}^{l}_{\text{student}}$ to align with the teacher’s $\mathbf{h}^{l}_{\text{teacher}}$, we are able to transfer the reasoning capability from explicit CoT to implicit CoT.
The Distilled Token.
Rather than aligning with all tokens in the query sentence, we select a distillation token for alignment. Inspired by the recent observations (Orgad et al., 2025) that the hidden activations of the token intermediately preceding the answer, i.e., the colon (“:”) in the answer prompt “The answer is:” (as shown in Figure 2), encodes essential reasoning information. We select this token’s hidden activations, h, for distillation.
Loss Function.
As a result, we formulate a loss function that aligns the teacher’s and student’s hidden activations across all layers at the selected distillation token for the student’s implicit CoT learning. To ensure a one-way flow of knowledge, we apply a stop-gradient operation on $\mathbf{h}^{l}_{\text{teacher}}$, only allowing the teacher to influence the student:
@@ 3.4 Self-Distillation
where $M$ indicates the number of layers in the LLM, sg denotes the stop-gradient operation, and $\textbf{h}^{l}$ is the hidden activations of the LLM’s $l$-th layer for the token position corresponding to the colon “:” in our design.
@@ 3.5 Training and Inference
The continuous thoughts are generated dynamically during training, as they are not known beforehand. To achieve this, we decode them step by step, with a cache storing previous keys and values to maintain efficiency. When applying a distance metric between two hidden activations, we observed significant norm variations across layers Deng et al. (2023); Cheng and Durme (2024). To address this, we normalize each layer’s hidden activations by dividing them by the standard deviation of the teacher’s corresponding hidden activations within the current batch.
For the distillation task, we adopt the same model for both the teacher and student roles for two primary reasons. (1) Reference Learning: The model must first learn to perform explicit CoT reasoning before it can effectively compress and transfer this capability into continuous space as implicit CoT. (2) Training Efficiency: While it is feasible to train separate teacher and student models—as explored in Section 4.4—this setup introduces additional complexity. The teacher must be pre-trained, and maintaining two distinct models during training doubles memory consumption.
For training data, we exclude the final CoT step—the step responsible for generating the final answer—because including this step could allow the teacher’s hidden activations to take a shortcut. Specifically, the model might directly copy the result from the last CoT step to the token responsible for generating the exact answer token, bypassing the reasoning process. This behavior would undermine the quality of the target hidden activations, as they would no longer fully encode the reasoning patterns. The ablation results demonstrating the impact of this exclusion are presented in Table 2.
@@
Figure 3: Results on five datasets (Top: GPT-2, Bottom: LLaMa3.2-1b-Instruct). CODI consistently outperforms all previous implicit CoT methods by a substantial margin. When using GPT-2, CODI even matches the performance of CoT-SFT on the in-domain GSM8k and GSM8k-NL datasets.